The product of two consecutive integers is 71 more than their sum. Find the integers.
Answer by richwmiller(17219)

n*(n+1)=n+n+1+71
n^2+n=2n+72
n^2-n-72=0-
(n-9)(n+8)=0
n=9 (and 10) and n= -8 (and -7)
check
9*10=19+71
90=90
-8*-7=-8-8+1+71
56=72-16
56=56